OpenStudy (anonymous):

If g(x) = x2 - 2x - 8, solve g(x) ≤ 0

OpenStudy (anonymous):

Hmm... Wanna learn how to do this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We can make it 'less scary' by changing g(x) to y.

OpenStudy (anonymous):

so, y = x^2 -2x -8 We want y to equal less than or equal to 0.

OpenStudy (anonymous):

Do you know how to factor this?

OpenStudy (anonymous):

As in how to factor y = x2 - 2x - 8 ?

OpenStudy (anonymous):

(x+2)(x-4)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Perfect! Now, next question: What shape is this equation?

OpenStudy (anonymous):

idk :(

OpenStudy (anonymous):

Not a problem! Just remember: if you have x^2 in your equation, what will it be? A line? A circle? A square? A parabola? A triangle?

OpenStudy (anonymous):

a parabola

OpenStudy (anonymous):

{x|-2 ≤ x ≤ 4}; [-2, 4]

OpenStudy (anonymous):

perfect! Now, is our parabola shaped more like a U (opening up) or a ^ (opening down)?

OpenStudy (anonymous):

And yeah, that is our answer, but do you know *why*?

OpenStudy (anonymous):

Hint: to remember if a parabola is going up or down, look at the sign in front of the x^2

OpenStudy (anonymous):

down because of the nagative???

OpenStudy (anonymous):

g(x) = x2 - 2x - 8

OpenStudy (anonymous):

That was our original question.

OpenStudy (anonymous):

So is there a positive or a negative in front of the x squared?

OpenStudy (anonymous):

so its facing up

OpenStudy (anonymous):

Perfect! And when you factored as (x-4)(x+2), where does that mean we place our points on the graph?

OpenStudy (anonymous):

Hint: it's (x-4)(x+2)=0 where the y is 0

OpenStudy (anonymous):

x=4 so 4,0 x=-2 so -2,0???

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

Now, here's where we solve: where on the graph is the line below y = 0? Here's a picture of it: http://www.wolframalpha.com/input/?i=g%28x%29+%3D+x^2+-+2x+-+8

OpenStudy (anonymous):

(That's why I asked for the direction of the graph. If it was upside down, our answer would be totally different)

OpenStudy (anonymous):

thank you so much . i apreciate that you helped me understand

OpenStudy (anonymous):

No problemo! Feel free to 'fan' me if you want to find me easily for future help! I like walking folks through to let them find answers on their own like this. :P

OpenStudy (anonymous):

Yay! A new fan.